OpenStudy (anonymous):

Using the given zero, find one other zero of f(x). Explain the process you used to find your solution. 1 - 6i is a zero of f(x) = x4 - 2x3 + 38x2 - 2x + 37.

OpenStudy (dumbcow):

given \[x = 1-6i\] then \[x-1 = -6i\] \[(x-1)^{2} = (-6i)^{2}\] \[x^{2} -2x+1 = -36\] \[x^{2} -2x +37 = 0\] Now this is one of the factors of polynomial...to get other factor divide f(x) by this factor x^2 + 1 ____________________________ x^2 -2x +37 | x^4 -2x^3 +38x^2 -2x+37 -(x^4 -2x^3+37x^2) ------------------ x^2 -2x +37 -(x^2 -2x +37) -------------- 0 to get other zeros, set factor equal to 0 \[x^{2} +1 = 0\] \[x^{2} = -1\] \[x = \pm i\]